Hannah's mother is 4 times as old as Hannah. After 16 years, her mother will be 2 times as old as
Hannah. Hannah is now .....years old and her mother is.... years old now.

Answer by ankor@dixie-net.com(22740)

Hannah's mother is 4 times as old as Hannah.
let h = Hannah's age
then
4h = her Mom's age
:
After 16 years, her mother will be 2 times as old as Hannah.
4h + 16 = 2(h+16)
4h + 16 = 2h + 32
4h - 2h = 32 - 16
2h = 16
h = 8 yrs is Hannah's age,
then
4(8) = 32 is Mom's age
:
you can confirm this in the statement
"After 16 years, her mother will be 2 times as old as Hannah"